# Linear Discriminant Regularized Regression

### v1

@@ 4 Simulation study
We generate the data from the following mean-shift scenario. A similar setting is also used in Witten and Tibshirani (2011). Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $\mu_{\ell,j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0. Regarding the within-class covariance matrix, we set $\Sigma_{w}=\sigma^{2}W$ where $W$ is generated by independently sampling its diagonal elements from Uniform(1,3), and setting its off-diagonal elements as $W_{ij}=\sqrt{W_{ii}W_{jj}}\rho^{|i-j|}$, for all $i\neq j$. The quantity $\sigma>0$ controls the overall noise level while the coefficient $\rho\in[0,1]$ controls the within-class correlation among the features. We generate the class probabilities $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ as follows. For a given $\alpha\geq 0$, we set $\pi_{\ell}=\nu_{\ell}^{\alpha}/(\sum_{i=1}^{L}\nu_{i}^{\alpha})$ for all $\ell\in[L]$. Here $\nu_{1},\ldots,\nu_{L}$ are independent draws from the Uniform$(0,1)$. The quantity $\alpha$ controls the imbalance of each class size: a larger $\alpha$ corresponds to more imbalanced class sizes. At the other extreme, $\alpha=0$ corresponds to the balanced case $\pi=(1/L)1_{L}$. In the following, we examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{100,300,500,700,900\}$; the feature dimension $p\in\{100,300,500,700,900\}$; the number of classes $L\in\{2,5,8,11,14\}$; the correlation coefficient $\rho\in\{0,0.2,0.4,0.6,0.8\}$; the noise parameter $\sigma\in\{0.6,0.9,1.2,1.5,1.8\}$; the imbalance of the class probabilities $\alpha\in\{0,0.4,0.8,1.2,1.6\}$.
@@
For all settings, we fix $n=300$, $p=500$, $L=5$, $\rho=0.6$, $\sigma=1$ and $\alpha=0$ when they are not varied. We use $500$ test data points to compute the misclassification errors of each procedure, and for each setting, we consider $50$ repetitions and report the averaged misclassification errors in Figure 1.
@@
From Figure 1, it is clear that our procedures consistently outperform other methods across all settings. We observe that the performance of our algorithms gets improved as $n$ increases, $L$ decreases and $\sigma$ decreases. Moreover, neither the imbalance of the class probabilities nor the ambient dimension $p$ seems to affect the misclassification error of our algorithms. These findings are in line with our theory in Section 3. It is a bit surprising to see that the classification error of our benchmark decreases as the correlation coefficient $\rho$ keeps increasing after $0.4$. Nevertheless, both our algorithms and RDA adapt to this situation. Finally, we find that both LDRR and LDRR-F as well as their lasso and elastic net variants have very comparable performance in all scenarios. This is as expected as the data is simulated according to model (1.1) and (1.3), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
@@ 5 Real data analysis
In this section we evaluate the performance of our approach and compare it with other algorithms on three high-dimensional biological data sets that are commonly used in the literature for discriminant analysis.
@@
The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001). The SRBCT dataset (Khan et al., 2001) has $n=83$ samples and consists of microarray analyses of childhood cancer tumors, more specifically, small round blue cell tumors (SRBCT). These tumors are categorized into $L=4$ classes: Burkitt lymphoma, Ewing sarcoma, neuroblastoma, and rhabdomyosarcoma. Each sample has $p=2,308$ gene expressions as features. The third dataset has $n=62$ samples and consists of $L=3$ distinct lymphoma subtypes. These lymphoma are categorized into diffuse large B-cell lymphoma, follicular lymphoma and chronic lymphocytic leukemia. Each sample has $p=4026$ gene expressions as its features. We used the preprocessed Lymphoma data in Dettling (2004).
@@
For all three data sets, the features are centered to zero mean. Each dataset is randomly divided with 75% of samples as the training data and 25% as the test data. Such process was repeated 10 times and the averaged misclassification errors of each algorithm were reported in Table 1.
Table 1: The averaged misclassification errors (in percentage).
LDRR PenalizedLDA PAMR RDA
Ramaswamy 14.4 37.6 28.8 15.4
SRBCT 1.0 5.2 1.4 1.0
Lymphoma 0.0 0.6 0.6 0.6
@@
Given the similar performance between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification. From Table 1, we see that our procedure LDRR-F has the smallest misclassification errors across all three datasets. Furthermore, in Figure 2, we show that LDRR-F with $K=2$ in (2.15) can be used for visualization in the 2D space of the first two discriminant directions.
@@
Figure 2: Visualization of Ramaswamy, SRBCT and Lymphoma data points in the space of the first two discriminant vectors.
@@ Acknowledgements
Appendix A Proofs
@@
$\Delta={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}$
@@
$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (A.1)
for some absolute constant $C>0$. Lemma 15 in Appendix C ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1. Also note that, on $\mathcal{E}_{\pi}$ and under the condition $L\log(n)\leq n$,
$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (A.2)
@@
A.1 Proof of the equivalence between (1.4) and (2.13)
@@
Recall that $B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}$. We find that
$B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}$ $=\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}$
@@
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$
$\qquad+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}).$
Write the singular value decomposition of $\Sigma_{w}^{-1/2}{\mathbb{M}}$ as $U\Lambda V^{\top}$ with $U\in\mathbb{R}^{p\times r}$ containing the first $r$ singular vectors and $r={\rm rank}(\Sigma_{w}^{-1/2}{\mathbb{M}})$. It follows that the above equals to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{\mathbb{M}}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{\mathbb{M}}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{\mathbb{M}}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{\mathbb{M}}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
@@
A.2 Proof of Theorem 2
@@
We work on the intersection of the event

### v2

@@ 4.1 Sparse scenarios
We generate the data from the following mean-shift scenario. A similar setting is also used in Witten and Tibshirani (2011). Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $[\mu_{\ell}]_{j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0. Regarding the within-class covariance matrix, we set $\Sigma_{w}=\sigma^{2}W$ where $W$ is generated by independently sampling its diagonal elements from Uniform(1,3), and setting its off-diagonal elements as $W_{ij}=\sqrt{W_{ii}W_{jj}}\rho^{|i-j|}$, for all $i\neq j$. The quantity $\sigma>0$ controls the overall noise level while the coefficient $\rho\in[0,1]$ controls the within-class correlation among the features. We generate the class probabilities $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ as follows. For a given $\alpha\geq 0$, we set $\pi_{\ell}=\nu_{\ell}^{\alpha}/(\sum_{i=1}^{L}\nu_{i}^{\alpha})$ for all $\ell\in[L]$. Here $\nu_{1},\ldots,\nu_{L}$ are independent draws from the Uniform$(0,1)$. The quantity $\alpha$ controls the imbalance of each class size: a larger $\alpha$ corresponds to more imbalanced class sizes. At the other extreme, $\alpha=0$ corresponds to the balanced case $\pi=(1/L)1_{L}$. In the following, we examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{100,300,500,700,900\}$; the feature dimension $p\in\{100,300,500,700,900\}$; the number of classes $L\in\{2,5,8,11,14\}$; the correlation coefficient $\rho\in\{0,0.2,0.4,0.6,0.8\}$; the noise parameter $\sigma\in\{0.6,0.9,1.2,1.5,1.8\}$; the imbalance of the class probabilities $\alpha\in\{0,0.4,0.8,1.2,1.6\}$.
@@
For all settings, we fix $n=300$, $p=500$, $L=5$, $\rho=0.6$, $\sigma=1$ and $\alpha=0$ when they are not varied. We use $500$ test data points to compute the misclassification errors of each procedure, and for each setting, we consider $50$ repetitions and report their averaged misclassification errors in Figure 1.
@@
From Figure 1, it is clear that our procedures consistently outperform other methods across all settings. We observe that the performance of our algorithms improves as $n$ increases, $L$ decreases and $\sigma$ decreases. Moreover, neither the imbalance of the class probabilities nor the ambient dimension $p$ seems to affect the misclassification error of our algorithms. These findings are in line with our theory in Section 3.3. It is a bit surprising to see that the classification error of our benchmark decreases as the correlation coefficient $\rho$ keeps increasing after $0.4$. Nevertheless, both our algorithms and RDA adapt to this situation. Finally, we find that both LDRR and LDRR-F as well as their lasso and elastic net variants have very comparable performance in all scenarios. This is as expected as the data is simulated according to model (1.1), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
@@
4.2 Low-rank scenarios
In this section we evaluate our proposed procedure when the rank of the conditional mean matrix $M$ is low and hence $B$ is also low-rank. We consider a general within-class covariance matrix $\Sigma_{w}$ here, and defer to Appendix A for the factor model mentioned in Remark 1 where $\Sigma_{w}$ has approximately low-rank.
Specifically, let the within-class covariance matrix $[\Sigma_{w}]_{ij}=\rho^{|i-j|}$ with $\rho=0.6$. To generate $M$, for some scalar $\eta>0$ and some randomly generated orthogonal matrix $A\in\mathbb{R}^{p\times r}$ with $A^{\top}A=\mathbf{I}_{r}$, we set $M=\eta A\alpha$ with entries of $\alpha\in\mathbb{R}^{r\times L}$ generated i.i.d. from $N(0,32/r)$. The quantity $\eta$ controls the magnitude of separation between the conditional means.
We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.3,0.6,0.9,1.2,1.5\}$. For each setting, we fix $n=1000,L=10,r=3,p=100$ and $\eta=1$ when they are not varied. We set $n=500$ when we vary $\eta$. Figure 2 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
From Figure 2 we see that the proposed approaches of using the reduced-rank penalty generally outperform the other methods. As $L$ gets larger with $r$ fixed, the benefit of using reduced-rank penalty becomes more visible, in line with 9. When the dimension ($p$) increases and gets closer to the sample size, the performance of LDRR(RR) deteriorates quickly while LDRR(RR+L2) still performs the best. For relative large $p$, incorporating the ridge penalty is beneficial as it improves the estimation of $B$ when the sample covariance matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ becomes ill-conditioned.
Figure 2: The averaged misclassification errors of each algorithm.
@@ 5 Real data analysis
In this section we evaluate the performance of our approach and compare it with other algorithms on three high-dimensional biological data sets and one low-dimensional football dataset.
@@
The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001). The SRBCT dataset (Khan et al., 2001) has $n=83$ samples and consists of microarray analyses of childhood cancer tumors, more specifically, small round blue cell tumors (SRBCT). These tumors are categorized into $L=4$ classes: Burkitt lymphoma, Ewing sarcoma, neuroblastoma, and rhabdomyosarcoma. Each sample has $p=2,308$ gene expressions as features. The third dataset has $n=62$ samples and consists of $L=3$ distinct lymphoma subtypes. These lymphoma are categorized into diffuse large B-cell lymphoma, follicular lymphoma and chronic lymphocytic leukemia. Each sample has $p=4026$ gene expressions as its features. We used the preprocessed Lymphoma data in Dettling (2004). The Football dataset from kaggle [^1] https://www.kaggle.com/datasets/vivovinco/20222023-football-player-stats/data contains 2022 – 2023 football player stats per 90 minutes. We excluded categorical data such as the player’s nationality, name, team name, and league from our prediction. We used age, year of birth and performance data $(p=118)$ to predict the $(n=2689)$ football player’s position $(L=10)$.
@@
For all four data sets, the features are centered to have zero mean. Each dataset is randomly divided with 75% of the samples to serve as the training data and the remaining 25% serves as the test data. This randomization was repeated 50 times and the averaged misclassification errors of each algorithm are reported in Table 1.
Table 1: The averaged misclassification errors (in percentage) and standard errors
LDRR-F(L1+L2) LDRR-F(L12+L2) LDRR-F(RR+L2) PenalizedLDA PAMR RDA
Ramaswamy 30.7(0.051) 14.4(0.054) 22.6(0.048) 40.0(0.058) 30.4(0.058) 16.7(0.05)
SRBCT 0.7(0.022) 0.3(0.011) 0.1(0.007) 4.4(0.052) 1.0(0.019) 0.0(0.0)
Lymphoma 0.3(0.012) 0.5(0.017) 0.4(0.015) 1.9(0.034) 1.3(0.025) 1.3(0.025)
Football 31.5(0.013) 32.2(0.011) 30.5(0.008) 52.6(0.008) 43.8(0.016) 31.9(0.014)
@@
Given the similar performance in Section 4 between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty (L1+L2) and the reduced rank plus ridge penalty (RR+L2) with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification. We also include LDRR-F that uses the group-lasso penalty plus the ridge penalty (L12+L2) for comparison. From Table 1, we see that our procedure LDRR-F is close to have the best performance across all datasets. Choice of the best penalty depends on the data structure. For instance, when the data has the group-sparsity structure such as in Ramaswamy data set, the classifier using the group-lasso penalty has better performance; when the sparsity structure varies across the response levels such as in Lymphoma, the classifier using the elastic-net penalty performs better; when the data meets the low rank structure, the benefit of the classifier using the reduced-rank estimator is apparent. For the Football dataset with $p=118$ and $L=10$, the selected rank is found to be $\widehat{r}=3$. Furthermore, in Figure 3, we show that LDRR-F with $K=2$ in (2.16) can be used for visualization in the 2D space of the first two discriminant directions. The algorithm RDA has comparable performance in all settings, but it cannot be used for low-dimensional representation / visualization.
@@
Figure 3: Visualization of Ramaswamy, SRBCT, Lymphoma and Football data points in the space of the first two discriminant vectors.
@@ Acknowledgements
Appendix A Additional simulation study when the conditional mean matrix is low-rank and $\Sigma_{w}$ is approximately low-rank
In this setting, we consider the setting in Remark 1 and also compare with the method (PCLDA) in Bing and Wegkamp (2023). As mentioned in Remark 1, PCLDA is a two-step procedure similar to LDRR-F but with the first step done in a unsupervised way by using PCA. We generate the data according to model (1.1) with $M=\eta A\alpha$ as in Section 4.2 and $\Sigma_{w}=\eta^{2}A\Sigma_{z}A^{\top}+{\rm Cov}(W)$ where $[\Sigma_{z}]_{ij}=(0.6)^{|i-j|}$ for $i,j\in[r]$ and ${\rm Cov}(W)$ is generated the same as $\Sigma_{w}$ in Section 4.1 with $\rho=0.2$ and its diagonal elements from Uniform(0,1). The quantity $\eta$ represents the signal-to-noise ratio (SNR) between the low-dimensional signal $\eta AZ$ and the noise $W$. As pointed out in Bing and Wegkamp (2023), classification becomes easier as $\eta$ increases.
We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.5,1.0,2.0,4.0,6.0\}$. For all settings, we fix $n=1000,L=10,r=3,p=100$ and $\eta=2$ when they are not varied. We set $n=500$ when we vary $\eta$. Figure 4 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
From Figure 4 we have similar findings as in the low-rank setting in Section 4.2 that the proposed approaches of using the reduced rank penalty outperform other methods when $n$ is much larger than $p$. In the last setting when we vary the SNR $\eta$, it is interesting to notice that PCLDA, the procedure that reduces the dimension by PCA in its first step, has comparable performance to the LDRR-F(RR+L2) when $\eta$ is sufficiently large ($\eta\geq 2$ in these settings). This aligns with the minimax optimality of PCLDA derived in Bing and Wegkamp (2023) in the regime of moderate / large $\eta$. But when $\eta$ is small such as $\eta<2$, LDRR-F has benefit by reducing the dimension in a supervised way.
Figure 4: The averaged misclassification errors of each algorithm.
Appendix B Proofs
@@
$\Delta={M}^{\top}\Sigma_{w}^{-1}{M}$
@@
$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (B.1)
for some absolute constant $C>0$. Lemma 21 in Appendix D ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1. Also note that, on $\mathcal{E}_{\pi}$ and under the condition $L\log(n)\leq n$,
$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (B.2)
@@
B.1 Proof of the equivalence between (1.2) and (2.14)
@@
Recall that $B^{*}=\Sigma_{w}^{-1}{M}$. We find that
$B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}$ $=\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}$
@@
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$
$\qquad+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}).$
Write the singular value decomposition of $\Sigma_{w}^{-1/2}{M}$ as $U\Lambda V^{\top}$ with $U\in\mathbb{R}^{p\times r}$ containing the first $r$ singular vectors and $r={\rm rank}(\Sigma_{w}^{-1/2}{M})$. It follows that the above equals to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
@@
B.2 Proof of 2
For any $x\in\mathbb{R}^{p}$, write
For any $t_{1},\ldots,t_{L}\geq 0$, define the event
$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t_{\ell}\right\}.$ (B.3)
In the proof of 2, we use the notation $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$.
@@ B.2 Proof of 2
We condition on ${\bm{D}}$ throughout the proof. The law of $\mathbb{E}$ and $\mathbb{P}$ is with respect to the randomness of $(X,Y)$, which is independent of ${\bm{D}}$. By definition, the excess risk $\mathcal{R}(\widehat{g})$ equals to
$\sum_{k\in[L]}\pi_{k}\Bigl{\{}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k\}\mid Y=k\right]-\mathbb{E}\left[\mathbbm{1}\{g^{*}(X)\neq k\}\mid Y=k\right]\Bigr{\}}$
$=\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k,g^{*}(X)=k\}\mid Y=k\right]-\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)\neq k\}\mid Y=k\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=k\right]-\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)=\ell\}\mid Y=k\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\Bigl{\{}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=k\right]-\pi_{\ell}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=\ell\right]\Bigr{\}}.$
[…]
where, for all $k,\ell\in[L]$,
Fix any $t_{1},\ldots,t_{L}\geq 0$. Observe that the event $\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\cap\mathcal{E}_{t}$ implies
$0>2G^{(\ell|k)}(X)\overset{\mathcal{E}_{t}}{\geq}\widehat{G}_{k}(X)-\widehat{G}_{\ell}(X)-(t_{k}+t_{\ell})\geq-(t_{k}+t_{\ell}).$
By the basic inequality $1+z\leq\exp(z)$ for all $z\in\mathbb{R}$, we obtain that,
$\mathcal{R}(\widehat{g})$ $\leq\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$
$+{1\over 2}\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}(t_{k}+t_{\ell})\mathbb{E}\left[\mathbbm{1}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y=k\right\}\right].$
For the first term, we have
$\sum_{k\neq\ell}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$ $\leq(L-1)\sum_{k=1}^{L}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)=(L-1)\mathbb{P}(\mathcal{E}_{t}^{c}).$ (B.4)
Regarding the second term, it is bounded from above by
${1\over 2}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})\mathbb{P}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0\mid Y=k\right\}.$ (B.5)
Note that $G^{(\ell|k)}(X)\mid Y=k$ is normally distributed
$N\left(-{1\over 2}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})+\log\left(\pi_{\ell}/\pi_{k}\right),~{}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\right).$
By the mean-value theorem, the quantity in (B.5) is no greater than
${1\over 4}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})^{2}$ $\leq{1\over 2}\sum_{k=1}^{L}\pi_{k}\left[(L-1)t_{k}^{2}+\sum_{\ell\in[L]\setminus\{k\}}t_{\ell}^{2}\right]={1\over 2}\sum_{k=1}^{L}(L\pi_{k}+1)t_{k}^{2}.$
Together with (B.4) and $\pi_{k}\leq C/L$, the proof is complete. ∎
B.3 Proof of Theorem 3
Proof.
Pick any $\ell\in[L]$. Recall that
$|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|+|(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2|\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)|.$ (B.6)
Throughout the proof, we work on $\mathcal{E}_{\pi}$ in (B.1) which implies (B.2), an upper bound for the last term on the right of (B.6). Recall that $\mathbb{P}\{\mathcal{E}_{\pi}\}\geq 1-2n^{-2}$ and note that we consider both the randomness of $(X,Y)$ and that of ${\bm{D}}$ in this proof.
For the first term in (B.6), we invoke Lemma 22 with $k=\ell$ and $v=B^{*}_{\ell}$, use the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}=\Delta_{\ell\ell}$ and take a union bound over $\ell\in[L]$ to conclude that the event
$\mathcal{E}_{\mu}:=\left\{\left|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}\right|\leq\sqrt{3\Delta_{\ell\ell}}\sqrt{L\log(n)\over n},~{}\text{for all }\ell\in[L]\right\}$ (B.7)
has probability at least $1-2Ln^{-3}\geq 1-n^{-2}$. Regarding the second term in (B.6)
the Cauchy-Schwarz inequality gives
$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\Sigma_{w}^{-1/2}\mu_{\ell}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\ell\ell}}$ (B.8)
Finally, conditioning on $Y={\bm{e}}_{k}$, the Gaussian tail of $X\mid Y,{\bm{D}}$ gives that, for all $t\geq 0$,
$\mathbb{P}\left\{|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\geq|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+t\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}~{}\Big{|}~{}Y={\bm{e}}_{k},{\bm{D}}\right\}\leq 2e^{-t^{2}/2}.$
Choosing $t=2\sqrt{\log(n)}$, taking the union bound over $k\in[L]$ and unconditioning yield that for all $\ell\in[L]$,
$|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq\sum_{k=1}^{L}\pi_{k}|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\log(n)}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+4\log(n)}$ (B.9)
holds with probability at least $1-2Ln^{-2}$. The last step uses (B.8). Collecting (B.2), (B.7), (B.8) and (B.3) completes the proof. ∎
B.4 Proof of 4
Proof.
For any $\ell\in[L]$, let
${t_{\ell}\over C}=\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
Recall from the proof of Theorem 3 that $1-\mathbb{P}(\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi})\leq 2n^{-2}+n^{-2}=3n^{-2}$ and
$\sum_{\ell=1}^{L}\mathbb{P}\left\{\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\geq t_{\ell}\right\}\cap\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\right\}\leq{1\over n}.$
By the fact $\mathcal{R}(\widehat{g})\leq 1$ almost surely and invoking 2, we find that
$\mathbb{E}_{{\bm{D}}}[\mathcal{R}(\widehat{g})]$ $\leq\mathbb{E}_{{\bm{D}}}\left[\min\left\{C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]$
$\quad+1-\mathbb{P}_{{\bm{D}}}\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}$